Find the equation of the family of curve having the given slope, and the equation of the family which passes through the given point.

m= 2x⁴/x; (0,5)

Note: the slope m = dy/dx
1
Expert's answer
2021-08-31T23:56:04-0400

Considering the following equation of the slope

"\frac{dy}{dx}= \frac{2x^4}{x}= 2x^3"

Separating the variables

"dy = 2x^3 dx\\
\int dy = \int 2x^3 dx\\
y= \frac{2x^4}{4}+C\\
y= \frac{x^4}{2}+C\\"

Finding the value of C

"y= \frac{x^4}{2}+C\\
5= \frac{0^4}{2}+C\\
C=5"

The equation of the curve is given as

"y= \frac{x^4}{2}+5\\"